# Generalization or Memorization: Data Contamination and Trustworthy Evaluation for Large Language Models

### v2

@@ Abstract
Recent statements about the impressive capabilities of large language models (LLMs) are usually supported by evaluating on open-access benchmarks. Considering the vast size and wide-ranging sources of LLMs’ training data, it could explicitly or implicitly include test data, leading to LLMs being more susceptible to data contamination. However, due to the opacity of training data, the black-box access of models, and the rapid growth of synthetic training data, detecting and mitigating data contamination for LLMs faces significant challenges. In this paper, we propose CDD, which stands for Contamination Detection via output Distribution for LLMs. CDD necessitates only the sampled texts to detect data contamination, by identifying the peakedness of LLM’s output distribution. To mitigate the impact of data contamination in evaluation, we also present TED: Trustworthy Evaluation via output Distribution, based on the correction of LLM’s output distribution. To facilitate this study, we introduce two benchmarks, i.e., DetCon and ComiEval, for data contamination detection and contamination mitigation evaluation tasks. Extensive experimental results show that CDD achieves the average relative improvements of 21.8%-30.2% over other contamination detection approaches in terms of Accuracy, F1 Score, and AUC metrics, and can effectively detect contamination caused by the variants of test data. TED significantly mitigates performance improvements up to 66.9% attributed to data contamination across 24 settings and 21 contamination degrees. In real-world applications, we reveal that ChatGPT exhibits a high potential to suffer from data contamination on HumanEval benchmark.  [^1] https://github.com/YihongDong/CDD-TED4LLMs
@@
Yihong Dong, Xue Jiang, Huanyu Liu, Zhi Jin, and Ge Li Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education; School of Computer Science, Peking University, Beijing, China {dongyh, jiangxue}@stu.pku.edu.cn, {zhijin, lige}@pku.edu.cn
@@ 1 Introduction
Although acknowledged the significance, data contamination detection and trustworthy evaluation for LLMs still persist as open and challenging issues Yang et al. (2023); Huang et al. (2023). The difficulties of data contamination detection can be essentially attributed to three factors: 1) Opaque Training Data. It is usually non-public and comprehensive, while continuously evolving for new LLMs. 2) Black Box Models. The parameters and output probabilities of LLMs may not be available, such as ChatGPT and GPT-4 OpenAI (2023). 3) Proliferation of Synthetic Data. It could introduce the variants [^2] These variants may include, but are not limited to, translations into other languages, additions of explanations or intermediate processes, and provisions of alternate solutions. of test data to training data. Further, the evaluation to mitigate the impact of data contamination has hardly been studied.
@@
We construct two new datasets, i.e., DetCon and ComiEval, for data contamination detection and contamination mitigation evaluation tasks, respectively. Experimental results demonstrate that CDD achieves state-of-the-art (SOTA) performance and is also suitable for identifying variant contamination, i.e., existing the variants of test data in training data. TED successfully mitigates the impact of data contamination in evaluation across various scenarios. Furthermore, we also provide strong evidence that ChatGPT suffers from data contamination on HumanEval dataset.
@@ 3.2 CDD for Data Contamination Detection
However, $\rho^{\prime}(d)$ assumes that test data must be leaked in its original form $\{x,y\}$, and does not take into account the possible contamination of the variant form, i.e., $\{x,\hat{y}\}$.
@@ 4 Experiment
Code Generation HumanEval {CodeLlama, CodeGen} StarCoder data 1 : {0, 0.1K, 1K, 10K} {1e-3, 2e-4, 4e-8} [0, 20] {Original, Variant333}
@@
Approach Not Need Prob. Not Need Param. Not Need Other LLM Consider Variant Contamination
@@ 4.1 Dataset
We simulate data contamination by training LLMs using benchmark data. To cover various scenarios of data contamination, we consider different settings, including two domain benchmarks leaked on four LLMs, two contamination form (i.e. original and variant leaked data), using three different learning rates during training, four mixing ratios of leaked data with other training data, and 21 degrees of contamination (i.e., occurrences). The detailed statistics can be found in Table 1. Due to the high cost of large-scale pre-training, we employ LoRA Hu et al. (2022) to fine-tune the base models on these various settings. On this basis, we construct the DetCon and ComiEval datasets.
@@
contains 2224 data contamination detection tasks, covering two domains (code generation and logical reasoning) and two contamination forms (original and variant), which need to detect whether a specific LLM has contamination on a particular data. We randomly select the data from the leaked dataset and the LLM from the settings in Table 1, where occurrence 0 refers to ‘uncontaminated’ and the others denote ‘contaminated’.
@@
Figure 5: The effectiveness of CDD for data contamination detection in original and variant contamination forms.
@@ 4.2 Data Contamination Detection
We also evaluate the performance of CDD and other contamination detection approaches in identifying original and variant forms of data contamination, as shown in Figure 5. In the cases of original contamination, as the degree of contamination increases, the detection effectiveness across all approaches improves. CDD outperforms the other approaches at lower contamination degrees, which are more challenging to detect. In contrast, in the cases of variant contamination, CDD alone maintains robust performance, whereas the other approaches encounter significant limitations.
@@ 5 Related Work
Some work Pan et al. (2020); Zhou et al. (2023); Jacovi et al. (2023); Dodge et al. (2021) exposed the serious consequences of data contamination and urged attention to this problem. However, most currently released LLMs did not open their pre-training corpus, which poses a new challenge for data contamination detection. Recent work tried to detect contamination without access to the pre-training corpus Oren et al. (2023); Deng et al. (2023); Golchin et al. (2023). Min-kProb Shi et al. (2023) calculated the average of the ksmallest probabilities of generated tokens and considered it as contaminated if it exceeded a certain threshold. The work Li (2023) assumed that data leaked into the training set tends to exhibit lower perplexity and utilizes perplexity analysis for detection. However, they often require other model outputs (e.g. probability) in addition to text, presenting challenges in detecting closed-source LLMs like ChatGPT, and they ignore the potential contamination from variants of test data.
@@
Our work requires only sampled texts to detect LLM’s data contamination via output distribution and considers the potential variant contamination.

### v3

@@ Abstract
Recent statements about the impressive capabilities of large language models (LLMs) are usually supported by evaluating on open-access benchmarks. Considering the vast size and wide-ranging sources of LLMs’ training data, it could explicitly or implicitly include test data, leading to LLMs being more susceptible to data contamination. However, due to the opacity of training data, the black-box access of models, and the rapid growth of synthetic training data, detecting and mitigating data contamination for LLMs faces significant challenges. In this paper, we propose CDD, which stands for Contamination Detection via output Distribution for LLMs. CDD necessitates only the sampled texts to detect data contamination, by identifying the peakedness of LLM’s output distribution. To mitigate the impact of data contamination in evaluation, we also present TED: Trustworthy Evaluation via output Distribution, based on the correction of LLM’s output distribution. To facilitate this study, we introduce two benchmarks, i.e., DetCon and ComiEval, for data contamination detection and contamination mitigation evaluation tasks. Extensive experimental results show that CDD achieves the average relative improvements of 21.8%-30.2% over other contamination detection approaches in terms of Accuracy, F1 Score, and AUC metrics, and can effectively detect implicit contamination. TED substantially mitigates performance improvements up to 66.9% attributed to data contamination across various contamination setups. In real-world applications, we reveal that ChatGPT exhibits a high potential to suffer from data contamination on HumanEval benchmark. [^1] https://github.com/YihongDong/CDD-TED4LLMs
@@
Yihong Dong, Xue Jiang, Huanyu Liu, Zhi Jin, Bin Gu, Mengfei Yang, and Ge Li [^1] Corresponding author. Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education; School of Computer Science, Peking University, Beijing, China Beijing Institute of Control Engineering, China Academy of Space Technology {dongyh, jiangxue}@stu.pku.edu.cn, {zhijin, lige}@pku.edu.cn
@@ 1 Introduction
Although acknowledged the significance, data contamination detection and trustworthy evaluation for LLMs still persist as open and challenging issues Yang et al. (2023); Huang et al. (2023). The difficulties of data contamination detection can be essentially attributed to three factors: 1) Opaque Training Data. It is usually non-public and comprehensive, while continuously evolving for new LLMs. 2) Black Box Models. The parameters and output probabilities of LLMs may not be available, such as ChatGPT and GPT-4 OpenAI (2023). 3) Proliferation of Synthetic Data. It could implicitly introduce the variants [^2] These variants may include, but are not limited to, translations into other languages, additions of explanations or intermediate processes, and provisions of alternate solutions. of test data to training data. Further, the evaluation to mitigate the impact of data contamination has hardly been studied.
@@
We construct two new datasets, i.e., DetCon and ComiEval, for data contamination detection and contamination mitigation evaluation tasks, respectively. Experimental results demonstrate that CDD achieves state-of-the-art (SOTA) performance and is also suitable for identifying implicit contamination, i.e., existing the variants of test data in training data. TED successfully mitigates the impact of data contamination in evaluation across various scenarios. Furthermore, we also provide strong evidence that ChatGPT suffers from data contamination on HumanEval dataset.
@@ 3.2 CDD for Data Contamination Detection
However, $\rho^{\prime}(d)$ assumes that test data must be explicitly leaked in its original form $\{x,y\}$, and does not take into account the possible implicit contamination of the variant form, i.e., $\{x,\hat{y}\}$.
@@ 4 Experiment
Code Generation HumanEval {CodeLlama, CodeGen} StarCoder data 1 : {0, 0.1K, 1K, 10K} {1e-3, 2e-4, 4e-8} [0, 20] {Explicit, Implicit333}
@@
Approach Not Need Prob. Not Need Param. Not Need Other LLM Consider Implicit Contamination
@@ 4.1 Dataset
We simulate data contamination by training LLMs using benchmark data. To cover various scenarios of data contamination, we consider different settings, including two domain benchmarks leaked on four LLMs, two contamination form (i.e. explicit and implicit leaked data), using three different learning rates during training, four mixing ratios of leaked data with other training data, and 21 degrees of contamination (i.e., occurrences). The detailed statistics can be found in Table 1. Due to the high cost of large-scale pre-training, we employ LoRA Hu et al. (2022) to fine-tune the base models on these various settings. On this basis, we construct the DetCon and ComiEval datasets.
@@
contains 2224 data contamination detection tasks, covering two domains (code generation and logical reasoning) and two contamination forms (explicit and implicit), which need to detect whether a specific LLM has contamination on a particular data. We randomly select the data from the leaked dataset and the LLM from the settings in Table 1, where occurrence 0 refers to ‘uncontaminated’ and the others denote ‘contaminated’.
@@
Figure 5: The effectiveness of CDD for data contamination detection in explicit and implicit contamination forms.
@@ 4.2 Data Contamination Detection
We also evaluate the performance of CDD and other contamination detection approaches in identifying explicit and implicit forms of data contamination, as shown in Figure 5. In the cases of explicit contamination, as the degree of contamination increases, the detection effectiveness across all approaches improves. CDD outperforms the other approaches at lower contamination degrees, which are more challenging to detect. In contrast, in the cases of implicit contamination, CDD alone maintains robust performance, whereas the other approaches encounter significant limitations.
@@ 5 Related Work
Some work Pan et al. (2020); Zhou et al. (2023); Jacovi et al. (2023); Dodge et al. (2021) exposed the serious consequences of data contamination and urged attention to this problem. However, most currently released LLMs did not open their pre-training corpus, which poses a new challenge for data contamination detection. Recent work tried to detect contamination without access to the pre-training corpus Oren et al. (2023); Deng et al. (2023); Golchin et al. (2023). Min-kProb Shi et al. (2023) calculated the average of the ksmallest probabilities of generated tokens and considered it as contaminated if it exceeded a certain threshold. The work Li (2023) assumed that data leaked into the training set tends to exhibit lower perplexity and utilizes perplexity analysis for detection. However, they often require other model outputs (e.g. probability) in addition to text, presenting challenges in detecting closed-source LLMs like ChatGPT, and they ignore the potential implicit contamination from variants of test data.
@@
Our work requires only sampled texts to detect LLM’s data contamination via output distribution and considers the potential implicit contamination.
@@ 7 Limitations
8 Acknowledgments
This research is supported by the National Natural Science Foundation of China under Grant No.62192730, 62192733, 61832009, 62192731, 62072007, the Key Program of Hubei under Grant JD2023008.
@@ Appendix A Details of Dataset Construction
Appendix B Case Study
We display the first 10 samples and the sample with greedy search is marked in bold as shown below.